# Directional Sheaf Hypergraph Networks: Unifying Learning on Directed and Undirected Hypergraphs

### v1

@@ B.3 Extended results
Appendix C Experimental Setup
C.1 Hardware Details
@@
C.2 Selected Baselines
We compare our models against twelve state-of-the-art methods from the hypergraph learning literature. From the undirected hypergraph-learning literature we include HGNN [Feng et al., 2019], HNHN [Dong et al., 2020], UniGCNII [Huang and Yang, 2021], LEGCN [Yang et al., 2022], HyperND [Tudisco et al., 2021], AllDeepSets and AllSetTransformer [Chien et al., 2022], ED-HNN [Wang et al., 2023a], SheafHyperGNN [Duta et al., 2023] and PhenomNN [Wang et al., 2023b]. From the directed hypergraph-learning literature, we consider GeDi-HNN [Fiorini et al., 2024] and DHGNN [Ma et al., 2024] as our baselines. DHGNN was originally designed for link prediction on directed graph datasets and relies on a learnable embedding table to represent node features. In our evaluation, we report the model’s performance using both this original embedding approach and an alternative setup with explicit node features. We adapted their implementation of the approximate Laplacian operator as it was originally tailored for directed graphs, despite their definition in principle working also for directed hypergraphs. Among the baselines present in the literature, we also examined DHMConv [Zhao et al., 2024], which is introduced as a spatial convolution for directed hypergraphs. In practice, though, its implementation is designed for directed 2-uniform hypergraphs (i.e. standard graphs). Unlike spectral approaches such as DHGNN, which leverage a Laplacian construction, spatial methods such as DHMConv rely on edge-wise indexing mechanisms that are inherently tied to a graph structure and cannot be meaningfully applied to the directed hypergraph setting considered in our work, where each hyperedge can contain multiple nodes and are not restricted to pair-wise relations.
C.3 Hyperparameter tuning
For tuning all the models, we employ a Bayesian optimization method. All models are trained for up to 500 epochs with early stopping set to 200 epochs. We employ Adam [Kingma and Ba, 2017] for optimizing the model parameters with $\text{lr}\in\{0.02,\,0.01,\,0.005,\,0.001\}$, $\text{wd}\in\{0,\,5\times 10^{-5},5\times 10^{-4}\}$. For all the models, we adopt a dropout $\in\{0.1,0.2,\ldots,0.9\}$, and for each model that has a selectable number of layers for the final classifier we fix it to 2. For each baseline, we select a range of parameters consistent with those investigated in their respective original works.
@@ C.3 Hyperparameter tuning
C.4 Datasets Description
We follow the data splits proposed by Zhang et al. [2021] for the Telegram, Texas, Wisconsin, Cornell. For Chameleon and Squirrelm we adopt the splits proposed by Platonov et al. [2023]. For Roman-empire and Amazon-Ratings we adopt the splits proposed by Platonov et al. [2023] and adopt the splits of Chien et al. [2022] for the remaining ones. In all cases, the datasets are partitioned into 50% training, 25% validation, and 25% test samples. For the email-Enron and email-EU datasets and for all synthetic datasets, node attributes are not available. In these cases, we resort to structural features, representing each node by its degree. The statistics of the 12 real-world datasets as well as synthetic ones are provided in Table 4. The datasets used for the experiments are:
@@ C.4 Datasets Description
  • email-Enron, email-EU A corporate email communication network built from Enron’s message logs. Nodes correspond to email accounts and edges record sender interactions. As ground-truth labels are unavailable, we derive node classes via the Spinglass community detection method Reichardt and Bornholdt [2006].
@@
  • Roman-empire The dataset is based on the Roman Empire article from English Wikipedia, which was selected since it is one of the longest articles on Wikipedia and it follows the construction proposed by Platonov et al. [2023]. Each node in the graph corresponds to one (non-unique) word in the text.
  • Amazon-ratings The dataset, as proposed by Platonov et al. [2023], is based on the Amazon product co-purchasing network metadata dataset from SNAP Datasets Leskovec and Krevl [2014]. Nodes are products (books, music CDs, DVDs, VHS video tapes), and edges connect products that are frequently bought together.
  • Synthetic Introduced in Fiorini et al. [2024] by following the methodology adopted in Zhang et al. [2021], these datasets are built as follows: a vertex set $V$ is partitioned into $c$ equally sized classes $C_{1},\ldots,C_{c}$. For each class $C_{i}$, we sample $I_{i}$ intra-class hyperedges that are undirected. The cardinality of each hyperedge is drawn uniformly from $\{h_{\min},\ldots,h_{\max}\}$, and its nodes are sampled uniformly from $C_{i}$. For each ordered pair of distinct classes $(C_{i},C_{j})$ with $i<j$, we create $I_{o}$ inter-class directed hyperedges. For every such hyperedge $e$, the tail set $T(e)$ is sampled from $C_{i}$ and the head set $H(e)$ from $C_{j}$; the sizes $|T(e)|$ and $|H(e)|$ are drawn uniformly from $\{h_{\min},\ldots,h_{\max}\}$. This induces a directional flow from $C_{i}$ to $C_{j}$ only when $i<j$. Using this procedure, we generate three datasets with $n=500$ nodes, $c=5$ classes, $h_{\min}=3$, $h_{\max}=10$, $I_{i}=30$ intra-class hyperedges per class, and an increasing number of inter-class directed hyperedges $I_{o}\in\{10,30,50\}$.
Table 4: Statistics of the datasets used in our experiments. Reported are the number of nodes, features, hyperedges, and classes, as well as the average hyperedge size ($\lvert e\rvert$), the average node degree ($\lvert v\rvert$), and the clique-expansion (CE) homophily computed as in Wang et al. [2023a].
@@
C.5 Directed Hypergraph from a Directed Graph
Figure 5: Example of the creation of a directed hyperedge from the out-neighborhood of a node. Suppose we have a graph where node $n1$ connects to nodes $n2$, $n3$, and $n4$, so that $(n1,n2)$, $(n1,n3)$, and $(n1,n4)$ belong to $E$. The construction procedure yields a directed hyperedge $e_{1}$ with tail set $T(e_{1})=\{n1\}$ and head set $H(e_{1})=\{n2,n3,n4\}$.
@@ C.5 Directed Hypergraph from a Directed Graph
This formulation preserves the source-target semantics of the original graph by expressing them as a higher-order relation. Such hyperedges are often referred to as forward directed hyperedges [Gallo et al., 1993]. When every hyperedge is forward directed, the structure is a forward directed hypergraph, which is the case for all real-world datasets considered in this work.
Appendix D On previous proposals of the Sheaf Hypergraph Laplacian
In this section, we revisit the definition of the Sheaf Hypergraph Laplacian proposed in Duta et al. [2023], noting that it fails to satisfy basic spectral properties expected of a Laplacian operator, most notably positive semidefiniteness. This shortcoming motivates our formulation, which, as discussed in Section 3.4, constitutes (to our knowledge) the first definition of a Sheaf Hypergraph Laplacian that is fully consistent with the spectral requirements of a convolutional operator also in the undirected setting. The (called linear in the paper—the nonlinear one is, in essence, the Laplacian of a 2-uniform hypergraph) Laplacian of Duta et al. [2023] is defined as follows:

### v2

@@ B.3 Extended Results
Additionally, we evaluate our method on two real-world directed hypergraph dataset for molecular reaction reframed as a hyperedge classification task, results are provided in Table 4. These datasets are the result of the merging of data from different sources such as Kearnes et al. (2021); Reizman et al. (2016); Lugo-Martinez et al. (2021) and are built inspired by Restrepo (2024), which proposes a novel way of modeling molecular reactions through directed hypergraphs. Dataset-1 contains 100,523100{,}523 nodes and 50,01650{,}016 hyperedges, with a total of 10 classes. Dataset-2 contains 956 nodes and 3,0213{,}021 hyperedges to classify among 6 different classes. These datasets consist of inherently directional hyperedges as they contain the molecular reactions expressed as set of reagents (the tail set) and set of products (the head set) composing a molecular reaction. The nodes’ features are built based on Morgan Fingerprints (Rogers and Hahn, 2010), which are one of the most widely used molecular descriptors. We employ the F1-score metric since the data has an imbalanced amount of samples for each class as shown in Section D.6.
As shown in Table 4, DSHN consistently outperforms all competing methods from both the undirected and directed hypergraph learning literature. On Molecular-1, it achieves an F1-score of 82.32%, improving upon the strongest baseline, GeDi-HNN, by 1.98%. On Molecular-2, DSHN attains 89.09%, exceeding AllSetTransformer by a relative margin of 1.37%.
Table 4: Mean F1-score and standard deviation for hyperedge classification on two molecular reaction datasets (test F1-score $\pm$ std). The best score is shown in bold, and the second-best is underlined.
Method Molecular-1 Molecular-2
HGNN 69.38 $\pm$ 0.48 81.40 $\pm$ 2.68
HNHN 32.27 $\pm$ 1.30 45.69 $\pm$ 7.48
UniGCNII 72.00 $\pm$ 0.59 85.61 $\pm$ 2.63
LEGCN OOM 84.75 $\pm$ 2.68
HyperND 44.16 $\pm$ 1.27 82.86 $\pm$ 3.17
AllDeepSets 79.17 $\pm$ 0.53 85.78 $\pm$ 3.01
AllSetTransformer 79.24 $\pm$ 1.08 87.89 $\pm$ 2.87
ED-HNN 66.37 $\pm$ 2.62 87.05 $\pm$ 1.96
SheafHyperGNN 57.99 $\pm$ 2.75 80.25 $\pm$ 2.20
PhenomNN 47.71 $\pm$ 2.90 86.27 $\pm$ 2.40
GeDi-HNN 80.72 $\pm$ 0.78 85.64 $\pm$ 2.42
DHGNN OOM 85.93 $\pm$ 3.49
DSHN OOM 89.09 $\pm$ 3.08
DSHNLight 82.32 $\pm$ 0.56 86.52 $\pm$ 2.68
Appendix C Implementation details
C.1 Computational Complexity
Comparison between DSHN and DSHNLight
Table 5: DSHN vs DSHNLight– FLOPS and Step Time (in ms) Analysis Across Different Datasets (Mean $\pm$ Standard Deviation)
Avg FLOPs/epoch()($\downarrow$) Avg step time $(\downarrow)$
Dataset DSHN DSHNLight DSHN DSHNLight
Roman-empire OOM 12,898,147,996,391 $\pm$ 43,606 OOM 1050.30 $\pm$ 152.04
Amazon-ratings OOM 15,061,770,374,298 $\pm$ 0 OOM 1080.26 $\pm$ 159.91
Comparison between DSHN and other models
Fig. 5 reports the average test accuracy of five representative models under approximately the same parameter budget. The results indicate that model size alone does not explain the performance of DSHN. For instance, although SheafHyperGNN and ED-HNN have a comparable number of parameters, their accuracy is significantly lower, being these undirected methods. In contrast, DSHN achieves an improvement of about 8% over the strongest directed baselines, despite having the same number of parameters thanks to the expressive power associated to complex-valued and directional restriction maps. On the other hand, Table 6 shows a comparison of DSHNLight with other models. Both DSHNLight and SheafHyperGNN generally incur higher computational costs than traditional hypergraph neural networks, although achieving, at least in the case of DSHNLight, a substantially better accuracy. This overhead stems directly from the requirement to learn and apply restriction maps at every node-hyperedge incidence as detailed in the next paragraph.
Table 6: FLOPs and Parameter Count Across Datasets and Methods.
Dataset ED-HNN SheafHyperGNN DHGNN DSHNLight
FLOPs #Params FLOPs #Params FLOPs #Params FLOPs #Params
Cora $3\times 10^{9}$ 125,959125{,}959 $2\times 10^{11}$ 404,576404{,}576 $7\times 10^{10}$ 4,077,0514{,}077{,}051 $2.0\times 10^{11}$ 437,447437{,}447
Citeseer $5\times 10^{9}$ 271,174271{,}174 $3\times 10^{11}$ 985,440985{,}440 $3\times 10^{11}$ 12,751,31812{,}751{,}318 $3.1\times 10^{11}$ 1,018,5021{,}018{,}502
email-Enron $4\times 10^{8}$ 34,31134{,}311 $6\times 10^{8}$ 37,98437{,}984 $2\times 10^{7}$ 13,33413{,}334 $7.0\times 10^{8}$ 70,85570{,}855
email-EU $5\times 10^{9}$ 34,50634{,}506 $3\times 10^{10}$ 38,75238{,}752 $5\times 10^{8}$ 14,37214{,}372 $2.6\times 10^{10}$ 71,05071{,}050
Telegram $1\times 10^{9}$ 34,11634{,}116 $2\times 10^{9}$ 37,21637{,}216 $4\times 10^{7}$ 13,24113{,}241 $1.9\times 10^{9}$ 70,66070{,}660
Cornell $2\times 10^{8}$ 143,109143{,}109 $1\times 10^{9}$ 473,184473{,}184 $6\times 10^{8}$ 542,566542{,}566 $1.9\times 10^{9}$ 506,437506{,}437
Texas $2\times 10^{8}$ 143,109143{,}109 $1\times 10^{9}$ 473,184473{,}184 $6\times 10^{8}$ 542,566542{,}566 $1.9\times 10^{9}$ 506,437506{,}437
Wisconsin $3\times 10^{8}$ 143,109143{,}109 $2\times 10^{9}$ 473,184473{,}184 $1\times 10^{9}$ 658,370658{,}370 $3.0\times 10^{9}$ 506,437506{,}437
Chameleon $2\times 10^{9}$ 182,917182{,}917 $2\times 10^{10}$ 632,416632{,}416 $1\times 10^{10}$ 2,379,7832{,}379{,}783 $2.7\times 10^{10}$ 632,416632{,}416
Squirrel $9\times 10^{9}$ 167,813167{,}813 $1\times 10^{11}$ 572,000572{,}000 $7\times 10^{10}$ 4,924,1724{,}924{,}172 $1.3\times 10^{11}$ 605,253605{,}253
Roman-empire $2\times 10^{10}$ 54,16254{,}162 $\times 10^{13} 11$7,344117{,}344 $1\times 10^{12}$ 6,850,7786{,}850{,}778 $3\times 10^{13} 11$7,344117{,}344
Amazon-ratings $3\times 10^{10}$ 53,31753{,}317 $2\times 10^{13}$ 114,016114{,}016 $1\times 10^{12}$ 7,398,9337{,}398{,}933 $5\times 10^{13} 11$4,016114{,}016
Asymptotic Complexity
  1. 1.
    Linear Transformation. The feature transformation is defined as
    $\mathbf{X^{\prime}}=(\mathbf{I}_{n}\otimes\mathbf{W}_{1})\mathbf{X}\mathbf{W}_{2}$
    where $\mathbf{W}_{1}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{2}\in\mathbb{R}^{f\times f}$. The resulting complexity is $\mathcal{O}(n(d^{2}f+df^{2}))=\mathcal{O}(n(cd+cf))=\mathcal{O}(nc^{2})$, where $c=df$.
  2. 2.
    $\mathbf{Q}^{\vec{\mathcal{F}}}_{N}\mathbf{X^{\prime}}.$
    The sparsity pattern of $\mathbf{Q}^{\vec{\mathcal{F}}}_{N}$ comes directly from the incidence matrix: each hyperedge of size $|e|$ induces $|e|^{2}$ nonzero blocks through the outer product $\mathbf{B}^{(q)}(e,:)^{\dagger}\mathbf{B}^{(q)}(e,:)$. Summing across all hyperedges gives a total of $\mathsf{S}_{2}=\sum_{e\in\mathcal{E}}|e|^{2}=\mathcal{O}\big(m\bar{e}^{2})$ nonzero blocks, where $\bar{e}$ is the average hyperedge size [^2] One could also upper bound the $\mathsf{S}_{2}$ term with $\mathcal{O}\big(m{n}^{2})$, however, that approximation would be highly pessimistic, considering a fully-dense representation of the hypergraph, where each hyperedge connects all nodes.. Applying the Laplacian then requires $\mathcal{O}(m\bar{e}^{2}c)$ for diagonal maps and $\mathcal{O}(m\bar{e}^{2}dc)$ with non-diagonal maps,
  3. 3.
    Learning the Sheaf. Restriction maps are predicted as
    $\Phi(\mathbf{x}_{v},\mathbf{x}_{e})=\sigma\!\left(\mathbf{V}\bigl(\mathbf{x}_{v}\,\|\,\mathbf{x}_{e}\bigr)\right)$
    where $\mathbf{V}$ is a learnable transformation and $\sigma$ a nonlinearity. The resulting $f$-dimensional vector is then used as input to $V$ for every node-edge incidence. Indicating with $\bar{v}$ the average number of partecipations of a node to an hyperedge, the computational complexity is $\mathcal{O}(\bar{v}mc)$ in the diagonal case, and $\mathcal{O}(\bar{v}md^{2}c)$ in the non-diagonal case.
  4. 4.
    Constructing the Laplacian. In the hypergraph setting we assemble
    $\mathbf{Q}^{\vec{\mathcal{F}}}_{N}=\mathbf{D}_{V}^{-\frac{1}{2}}{\mathbf{B}^{(q)}}^{\dagger}\mathbf{D}_{E}^{-1}\mathbf{B}^{(q)}\mathbf{D}_{V}^{-\frac{1}{2}}.$
    The work naturally splits into two steps:
    1. (a)
      Degree normalization. This involves computing the node and hyperedge degree matrices, $\mathbf{D}_{V}^{-\tfrac{1}{2}}$ and $\mathbf{D}_{E}^{-1}$. For vertex degree normalization each node requires aggregating contributions from its incident hyperedges, giving $\mathcal{O}(m\bar{e}d)$ operations in the diagonal case and $\mathcal{O}(m\bar{e}d^{3})$ in the non-diagonal case (since each block is $d\times d$), to which it must be added the cost of inverting the block-diagonal matrices, adding to the complexity $\mathcal{O}\big(nd)$ in the diagonal case and $\mathcal{O}\big(nd^{3})$ in the non-diagonal case while since $\mathbf{D}_{E}^{-1}$ is obtained by expanding to matrix for the scalar hyperedge degrees $\delta_{e}$ this cost adds a trascurable term to the asymptotic complexity.
    2. (b)
      Sparse product. Forming the term
      $\mathbf{Q}^{\vec{\mathcal{F}}}_{N}=\mathbf{D}_{V}^{-\frac{1}{2}}{\mathbf{B}^{(q)}}^{\dagger}\mathbf{D}_{E}^{-1}\mathbf{B}^{(q)}\mathbf{D}_{V}^{-\frac{1}{2}}.$
      requires, for each hyperedge $e$, generating block interactions among all pairs of nodes it contains. This gives a total of $\mathsf{S}_{2}=\sum_{e\in\mathcal{E}}|e|^{2}$ block products. The cost is $\mathcal{O}(\mathsf{S}_{2}d)$ in the diagonal case and $\mathcal{O}(\mathsf{S}_{2}d^{3})$ in the non-diagonal case. Since the normalization terms $\mathbf{D}_{V}^{-\tfrac{1}{2}}$ are block-diagonal operations they do not contribute substantially in the overall complexity. Since $\mathsf{S}_{2}=\mathcal{O}\big(m\bar{e}^{2}\big)$, the dominant cost becomes $\mathcal{O}(m\bar{e}^{2}d)$ for diagonal maps and $\mathcal{O}(m\bar{e}^{2}d^{3})$ for non-diagonal maps.
By summing the overall contributions we get: $\mathcal{O}\!\left(n\,(c^{2}+d)\;+\;m\,(\,\bar{e}d+\bar{e}^{2}(d+c)+\bar{v}\,c\,)\right)$ in the diagonal case and $\mathcal{O}\!\left(n\,(c^{2}+d^{3})+m\,(\,\bar{e}d^{3}+\bar{e}^{2}(d^{3}+dc)+\bar{v}\,d^{2}c\,)\right)$ in the non-diagonal case.
Considerations on the Asymptotic Complexity
C.2 Architectural Choices
Layer Normalization
Each layer may optionally include layer normalization, with this choice considered a tunable hyperparameter, since it improves training stability and overall performance. Since the input signal to each convolutional layer is complex-valued, we adopt a complex normalization strategy as proposed in Trabelsi et al. (2018); Barrachina et al. (2023), where each complex feature is treated as a two-dimensional real vector $(\Re(x),\Im(x))$. Specifically, we compute the full $2\times 2$ covariance matrix:
$x_{o}=\gamma\tilde{x}+\beta,$
Residual Connections
Following observations from Bodnar et al. (2022), we optionally include residual connections in our convolutional layers, which we found to help the architecture in certain datasets. The use of residuals is treated as a tunable hyperparameter (see Section D.3). With this addition, a convolutional layer takes the form:
$\mathbf{X}_{t+1}=\sigma\left(\mathbf{Q}^{\vec{\mathcal{F}}}_{N}\,(\mathbf{I}_{n}\otimes\mathbf{W}_{1})\,\mathbf{X}_{t}\,\mathbf{W}_{2}+\mathbf{X}_{t}\right)\in\mathbb{C}^{nd\times f}.$
Activation Function
For the activation function, we adopt the complex ReLU commonly employed in related works (Zhang et al., 2021; Fiorini et al., 2023; 2024). It is defined as:
DSHNLight
Appendix D Experimental Setup
D.1 Hardware Details
@@
D.2 Selected Baselines
We compare our models against twelve state-of-the-art methods from the hypergraph learning literature. From the undirected hypergraph-learning literature we include HGNN (Feng et al., 2019), HNHN (Dong et al., 2020), UniGCNII (Huang and Yang, 2021), LEGCN (Yang et al., 2022), HyperND (Tudisco et al., 2021), AllDeepSets and AllSetTransformer (Chien et al., 2022), ED-HNN (Wang et al., 2023a), SheafHyperGNN (Duta et al., 2023) and PhenomNN (Wang et al., 2023b). From the directed hypergraph-learning literature, we consider GeDi-HNN (Fiorini et al., 2024) and DHGNN (Ma et al., 2024) as our baselines. DHGNN was originally designed for link prediction on directed graph datasets and relies on a learnable embedding table to represent node features. In our evaluation, we report the model’s performance using both this original embedding approach and an alternative setup with explicit node features.
D.3 Hyperparameter tuning
For tuning all the models, we employ a Bayesian optimization method. All models are trained for up to 500 epochs with early stopping set to 200 epochs. We employ Adam (Kingma and Ba, 2017) for optimizing the model parameters with $\text{lr}\in\{0.02,\,0.01,\,0.005,\,0.001\}$, $\text{wd}\in\{0,\,5\times 10^{-5},5\times 10^{-4}\}$. For all the models, we adopt a dropout $\in\{0.1,0.2,\ldots,0.9\}$, and for each model that has a selectable number of layers for the final classifier we fix it to 2. For each baseline, we select a range of parameters consistent with those investigated in their respective original works:
@@ D.3 Hyperparameter tuning
D.4 Datasets Description
We follow the data splits proposed by Zhang et al. (2021) for the Telegram, Texas, Wisconsin, Cornell. For Chameleon and Squirrel we adopt the splits proposed by Platonov et al. (2023). For Roman-empire and Amazon-Ratings we adopt the splits proposed by Platonov et al. (2023) and adopt the splits of Chien et al. (2022) for the remaining ones. In all cases, the datasets are partitioned into 50% training, 25% validation, and 25% test samples. For the email-Enron and email-EU datasets and for all synthetic datasets, node attributes are not available. In these cases, we resort to structural features, representing each node by its degree. The statistics of the 12 real-world datasets as well as synthetic ones are provided in Table 7. The datasets used for the experiments are:
@@ D.4 Datasets Description
  • email-Enron, email-EU A corporate email communication network built from Enron’s message logs. Nodes correspond to email accounts and edges record sender interactions. As ground-truth labels are unavailable, we derive node classes via the Spinglass community detection method Reichardt and Bornholdt (2006).
@@
  • Roman-empire The dataset is based on the Roman Empire article from English Wikipedia, which was selected since it is one of the longest articles on Wikipedia and it follows the construction proposed by Platonov et al. (2023). Each node in the graph corresponds to one (non-unique) word in the text.
  • Amazon-ratings The dataset, as proposed by Platonov et al. (2023), is based on the Amazon product co-purchasing network metadata dataset from SNAP Datasets Leskovec and Krevl (2014). Nodes are products (books, music CDs, DVDs, VHS video tapes), and edges connect products that are frequently bought together.
  • Synthetic Introduced in Fiorini et al. (2024) by following the methodology adopted in Zhang et al. (2021), these datasets are built as follows: a vertex set $V$ is partitioned into $c$ equally sized classes $C_{1},\ldots,C_{c}$. For each class $C_{i}$, we sample $I_{i}$ intra-class hyperedges that are undirected. The cardinality of each hyperedge is drawn uniformly from $\{h_{\min},\ldots,h_{\max}\}$, and its nodes are sampled uniformly from $C_{i}$. For each ordered pair of distinct classes $(C_{i},C_{j})$ with $i<j$, we create $I_{o}$ inter-class directed hyperedges. For every such hyperedge $e$, the tail set $T(e)$ is sampled from $C_{i}$ and the head set $H(e)$ from $C_{j}$; the sizes $|T(e)|$ and $|H(e)|$ are drawn uniformly from $\{h_{\min},\ldots,h_{\max}\}$.
Table 7: Statistics of the datasets used in our experiments. Reported are the number of nodes, features, hyperedges, and classes, as well as the average hyperedge size ($\lvert e\rvert$), the average node degree ($\lvert v\rvert$), and the clique-expansion (CE) homophily computed as in Wang et al. (2023a).
@@
D.5 Directed Hypergraph from a Directed Graph
@@
Figure 7: Example of the creation of a directed hyperedge from the out-neighborhood of a node. Suppose we have a graph where node $v_{1}$ connects to nodes $v_{2}$, $v_{3}$, and $v_{4}$, so that $(v_{1},v_{2})$, $(v_{1},v_{3})$, and $(v_{1},v_{4})$ belong to $E$. The construction procedure yields a directed hyperedge $e_{1}$ with tail set $T(e_{1})=\{v_{1}\}$ and head set $H(e_{1})=\{v_{2},v_{3},v_{4}\}$.
This formulation preserves the source-target semantics of the original graph by expressing them as a higher-order relation. Such hyperedges are often referred to as forward directed hyperedges (Gallo et al., 1993). When every hyperedge is forward directed, the structure is a forward directed hypergraph, which is the case for all real-world datasets considered in this work.
D.6 Hyperedge Classification for Molecular Reaction Type Prediction
In Table 8 and Table 9 we show the distribution of labels for the directed molecular reaction prediction datasets employed for the hyperedge classification task. As mentioned in Section B.3, we evaluate all models on additional real-world molecular prediction datasets (see Table 4) from the perspective of a hyperedge classification task. To do so, before feeding the output of the last convolutional layer to the classifier, we perform an aggregation (sum) of all node representations belonging to a given hyperedge. Specifically, if $\mathbf{X}_{\text{node}}$ is the final feature matrix of shape $N\times F$, where $N$ is the number of nodes and $F$ is the feature dimension, we can compute hyperedge-level representations using a (real-valued and binary) incidence matrix $\mathbf{H}$ of shape $N\times E$ as follows:
$\mathbf{H}^{\top}\mathbf{X}_{\text{node}}=\mathbf{X}_{\text{edge}}\in\mathbb{R}^{E\times F}.$
The resulting matrix $\mathbf{X}_{\text{edge}}$ is then fed to the classifier which will output a matrix of shape $E\times C$, where $C$ is the number of classes.
Class # Hyperedges Percentage (%)
0 15,151 30.29
1 11,896 23.78
2 5,662 11.32
3 909 1.82
4 672 1.34
5 8,237 16.47
6 4,614 9.23
7 811 1.62
8 1,834 3.67
9 230 0.46
Table 8: Label distribution for the Molecular-1 dataset.
Class # Hyperedges Percentage (%)
0 960 31.78
1 1,536 50.84
2 213 7.05
3 54 1.79
4 226 7.48
5 32 1.06
Table 9: Label distribution for the Molecular-2 dataset.
Appendix E On previous proposals of the Sheaf Hypergraph Laplacian
In this section, we revisit the definition of the Sheaf Hypergraph Laplacian proposed in Duta et al. (2023), noting that it fails to satisfy basic spectral properties expected of a Laplacian operator, most notably positive semidefiniteness. This shortcoming motivates our formulation, which, as discussed in Section 3.4, constitutes (to our knowledge) the first definition of a Sheaf Hypergraph Laplacian that is fully consistent with the spectral requirements of a convolutional operator also in the undirected setting. For comparison, we recall the (called linear in the paper—the nonlinear one is, in essence, the Laplacian of a 2-uniform hypergraph) Laplacian of Duta et al. (2023).